The mean annual cost of automobile insurance is R95. Assume that the population standard deviation is R14. What is the probability that a simple random sample of size 30 for automobile insurance policies will have a sample mean less than R90?


1
Expert's answer
2021-07-21T02:51:01-0400

Mean = R95, SD = R14 & N = 30

X = (R90 – R95) = -5

P(-5<X) =

Therefore,

Q = 14 /Square root of 30

=14/ 5.4772

= 2.5561

(90 – 95)/2.5561 = - 1.96

Probability is – 1.96